Find the distance of the point (0,2) from the straight line with slope 5 and passing through the point of intersection of x+2y=5 and x-3y=7. Find the equation of the straight line passing through the point of intersection of the straight lines x-3y+1=0 and 2x+5y-9=0 and having infinite slope.

  1. Find the distance of the point (0,2) from the straight line with slope 5 and passing through the point of intersection of x+2y=5 and x-3y=7.
  2. Find the equation of the straight line passing through the point of intersection of the straight lines x-3y+1=0 and 2x+5y-9=0 and having infinite slope.

Grade:12

2 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
the point of intersection of equations is(29/5,-2/5)
given slope is 5
So eqn of line is y-y1=m(x-x1)
y+2/5=5(x-29/5)
Similarly you can solve the next question
Thanks and regards
Shaik Aasif
askIITians faculty
SHAIK AASIF AHAMED
9 years ago
Hello student,
the point of intersection of equations is(29/5,-2/5)
given slope is 5
So eqn of line is y-y1=m(x-x1)
y+2/5=5(x-29/5)
5x-5y-31=0
Distance from point (0,2) to above straight line is\frac{5(0)-5(2)-31}{\sqrt{5^{2}+5^{2}}}
=21/5sqrt(2)=2.96
Similarly you can solvethe next question
Thanks and regards
Shaik Aasif
askIITians faculty